# Practical Unlearning for Large Language Models

### v1

@@ Abstract
While large language models (LLMs) have demonstrated impressive performance across various domains and tasks, their security issues have become increasingly severe. Machine unlearning (MU) has emerged as a promising solution to address these issues by aiming to remove the influence of undesired data on the target model without compromising its utility in other aspects. Traditional MU typically assumes full access to the original training data to preserve utility, which is difficult to achieve in LLM unlearning due to the vastness of the original training data. Therefore, existing LLM unlearning methods often assume access to data most affected by undesired data unlearning. However, this assumption underestimates the entanglement among various LLM capabilities and ignores data access limitations due to privacy concerns, copyright protection, and the potential rarity of the tasks. Moreover, these LLM unlearning methods do not sufficiently consider that unlearning requests in real-world scenarios are continuously emerging, which may lead to accumulated model utility loss that eventually becomes unacceptable. To overcome these challenges and achieve practical LLM unlearning, we propose the O3 framework. The O3 framework includes an Out-Of-Distribution (OOD) detector to measure the similarity between input and unlearning data, and an Orthogonal low-rank adapter (LoRA) for continuously unlearning requested data. The OOD detector is trained with a novel contrastive entropy loss and utilizes a local-global layer-aggregated scoring mechanism. The orthogonal LoRA achieves parameter disentanglement among continual unlearning requests. During inference, our O3 framework can smartly decide whether and to what extent to load the unlearning LoRA based on the OOD detector’s predictions. Notably, O3’s effectiveness does not rely on any retained data. Along with the LoRA design at the model level, this also makes O3 more computationally efficient. We conducted extensive experiments on O3 and state-of-the-art LLM unlearning methods across three tasks and seven datasets. The results indicate that O3 consistently achieves the best trade-off between unlearning effectiveness and utility preservation, especially when facing continuous unlearning requests.
11footnotetext: Equal contributions (ordered alphabetically). 2 Corresponding Author.
I Introduction
@@
Recently, bolstered by scaling laws [1], the size of language models has grown tremendously, demonstrating unexpectedly excellent performance across a variety of tasks [2]. However, concerns about large language models (LLMs) have also increased, particularly regarding safety alignment [3], hallucination output [4], and privacy violations [5]. To address these issues, more research focuses on applying different techniques [6] in LLMs to mitigate dangerous behaviors, eliminate incorrect knowledge, and remove private, toxic, or illegal data. Among these techniques, one of the most representative approaches is machine unlearning [7]. Current methods for LLM unlearning can be primarily categorized into parameter optimization [8, 9, 10, 11, 12, 13, 14], parameter merging [15, 16, 17, 18], and in-context learning [19, 20]. The parameter optimization methods involve directly fine-tuning the LLM, with the objective typically being to maximize the task loss on the data that needs to be unlearned or to minimize the random label loss on the unlearning data. Parameter merging for unlearning often requires identifying and localizing the model parameters related to the unlearning data and then making appropriate modifications to these parameters. Finally, in-context learning-based methods modify the LLM input prompts to make the LLM refuse to output content related to the unlearning data. In terms of unlearning effectiveness, parameter optimization and parameter merging methods are typically much more effective than in-context learning. However, parameter optimization and merging often poorly maintain the model utility outside the unlearning knowledge scope.
@@
Despite the above efforts, existing LLM unlearning studies still face the following challenges and struggle to be employed practically. First, in addition to the data that needs to be unlearned, existing unlearning methods also require a large dataset called the retained dataset to maintain the model utility. This retained dataset often consists of the original training dataset [7] or a portion of it. However, as LLMs are trained on massive datasets [2], assuming access to the complete training data to maintain model utility is typically unrealistic for LLM unlearning [21]. Moreover, as time goes on, the original training data of LLMs may become inaccessible due to data privacy protection, expired access authorization, and intellectual property protection [6]. If the retained dataset only contains incomplete training data distribution, the corresponding model utility for the missing parts will significantly decline during unlearning. Although some studies shrink the range of the retained data to the distribution most susceptible to unlearning, this distribution itself is hard to characterize and its corresponding data may be limited due to intrinsic rarity and privacy protection [22, 23, 24]. The second challenge is that LLM unlearning is often not a one-off operation but a continual process, as unlearning requests continuously emerge in the real world [21]. This is similar to LLM continual learning [25], except that LLM unlearning requires continuously unlearning outdated, incorrect, or user-requested information. As the number of unlearning operations increases, the aforementioned decline in model utility will also have a cumulative effect, meaning that the model’s general capabilities will significantly decrease [26, 27] over time. Existing LLM unlearning methods, however, only consider single operations and cannot perform effective continual unlearning.
@@
In this work, to achieve practical unlearning in LLMs, we propose the O3 framework. The O3 framework can balance unlearning effectiveness and model utility preservation in continuous scenarios without using any retained data. At a high level, the O3 framework mainly includes an Out-Of-Distribution (OOD) detection module to assess the similarity between input data and unlearning data, and an Orthogonal Low-rank adapter (LoRA) [28] for continuously unlearning requested data. Specifically, the OOD detector in O3 is trained with a novel contrastive entropy loss as its backbone. When supplemented with our local-global layer-aggregated scoring, this backbone can achieve truly unsupervised OOD detection, meaning that the available in-distribution training data is unlabeled and may be few-shot. The orthogonal LoRA in O3 enables the disentanglement of parameter space across different unlearning requests, ensuring that the unlearning effectiveness of different requests does not interfere with each other. The O3 framework can balance unlearning effectiveness and model utility because it smartly leverages the data similarity determined by the OOD detector to decide whether and to what extent to load the unlearning LoRA during inference. In summary, the main contributions of this work include:
  • We tackle important challenges for LLM unlearning in pratice: the continuous arrival of unlearning requests makes it hard to balance unlearning effectiveness and model utility preservation, and this is even harder when there is insufficient retained data.
@@
  • We propose a novel O3 framework that includes an OOD detector and an orthogonal unlearning LoRA. The OOD detector measures the similarity between input and unlearning data, allowing O3 to decide whether and to what extent to load the unlearning LoRA during inference. The orthogonal design of LoRA prevents interference among different unlearning requests, achieving better unlearning effectiveness in continuous scenarios.
@@
  • We conduct extensive experiments on three benchmark tasks that comprehensively test LLM’s discriminative, generative, and reasoning capabilities. The experiment results demonstrate that O3 consistently achieves the best trade-off between unlearning effectiveness and utility preservation without using any retained data, compared with many state-of-the-art baseline methods when facing continuous unlearning requests. Moreover, thanks to the design of LoRA and the elimination of the need for any retained data, O3 also provides improved computation efficiency.
II Background
II-A Machine Unlearning
Standard machine unlearning [7] ensures the “right to be forgotten” [29, 30] and allows individuals to request the deletion of their data from the machine learning service provider. Formally speaking, we assume the data that needs to be unlearned is termed as $\mathcal{D}^{\mathrm{U}}\!=\!\{(\bm{x}_{i},y_{i})\|\bm{x}_{i}\!\sim\!\mathcal{P}_{\mathcal{X}}^{\mathrm{U}},y_{i}\!\sim\!\mathcal{P}_{\mathcal{Y}}^{\mathrm{U}}\}_{i=1}^{N^{\mathrm{U}}}$ ($N^{\mathrm{U}}$ is the sample quantity). $\mathcal{P}_{\mathcal{X}}^{\mathrm{U}}$ and $\mathcal{P}_{\mathcal{Y}}^{\mathrm{U}}$ are the input and label marginal distributions of the unlearning set, respectively. Note that $\mathcal{D}^{\mathrm{U}}$ naturally comes from the original training dataset of the target model $M_{\Theta}$ that is parameterized on $\Theta$. Moreover, we denote the retained training dataset without $\mathcal{D}^{\mathrm{U}}$ as $\mathcal{D}^{\mathrm{R}}=\{(\bm{x}_{i},y_{i})\|\bm{x}_{i}\!\sim\!\mathcal{P}_{\mathcal{X}}^{\mathrm{R}},y_{i}\!\sim\!\mathcal{P}_{\mathcal{Y}}^{\mathrm{R}}\}_{i=1}^{N^{\mathrm{R}}}$. To achieve effective unlearning, the most direct way is to retrain another randomly initialized model from scratch with the retained dataset $\mathcal{D}^{\mathrm{R}}$. However, when the target model becomes larger and requires more training data, such training from scratch consumes a large volume of computation resources. Many approximation approaches have been proposed to reduce such high computational overhead, which we will discuss in Section VII. These approaches share the same objective, which is to remove the memorization about $\mathcal{D}^{\mathrm{U}}$ from the target model $M_{\Theta}$ while retaining the performance on $\mathcal{D}^{\mathrm{R}}$,
$\min_{\Theta^{\prime}}\mathbf{I}(M_{\bm{x}\in\mathcal{D}^{\mathrm{U}}}(\bm{x},\Theta);M^{\prime}_{\bm{x}\in\mathcal{D}^{\mathrm{U}}}(\bm{x},\Theta^{\prime}))$ (1)
$\max_{\Theta^{\prime}}\mathbf{I}(M_{\bm{x}\in\mathcal{D}^{\mathrm{R}}}(\bm{x},\Theta);M^{\prime}_{\bm{x}\in\mathcal{D}^{\mathrm{R}}}(\bm{x},\Theta^{\prime})),$

### v2

# On Large Language Model Continual Unlearning

@@ Abstract
While large language models have demonstrated impressive performance across various domains and tasks, their security issues have become increasingly severe. Machine unlearning has emerged as a representative approach for model safety and security by removing the influence of undesired data on the target model. However, these methods do not sufficiently consider that unlearning requests in real-world scenarios are continuously emerging, especially in the context of LLMs, which may lead to accumulated model utility loss that eventually becomes unacceptable. Moreover, existing LLM unlearning methods often ignore previous data access limitations due to privacy concerns and copyright protection. Without previous data, the utility preservation during unlearning is much harder. To overcome these challenges, we propose the O3 framework that includes an Orthogonal low-rank adapter (LoRA) for continually unlearning requested data and an Out-Of-Distribution (OOD) detector to measure the similarity between input and unlearning data. The orthogonal LoRA achieves parameter disentanglement among continual unlearning requests. The OOD detector is trained with a novel contrastive entropy loss and utilizes a glocal-aware scoring mechanism. During inference, our O3 framework can decide whether and to what extent to load the unlearning LoRA based on the OOD detector’s predicted similarity between the input and the unlearned knowledge. Notably, O3’s effectiveness does not rely on any retained data. We conducted extensive experiments on O3 and state-of-the-art LLM unlearning methods across three tasks and seven datasets. The results indicate that O3 consistently achieves the best unlearning effectiveness and utility preservation, especially when facing continuous unlearning requests. The source codes can be found at https://github.com/GCYZSL/O3-LLM-UNLEARNING.
1 Introduction
@@
Recently, bolstered by scaling laws (Kaplan et al., 2020), the size of language models has grown tremendously, demonstrating excellent performance across various tasks (Wang et al., 2024). However, concerns about large language models (LLMs) have also increased, particularly regarding how to eliminate undesirable data influence (e.g., privacy information (Pan et al., 2020)). To address this issue, machine unlearning (Bourtoule et al., 2021) is applied in LLMs to remove private, toxic, or illegal data. Current methods for LLM unlearning can be primarily categorized into parameter optimization (Chen & Yang, 2023; Eldan & Russinovich, 2023; Jia et al., 2024; Zhang et al., 2024; Meng et al., 2022; Li et al., 2024), and in-context unlearning (Thaker et al., 2024; Pawelczyk et al., 2024). The parameter optimization methods involve directly fine-tuning the LLM, with the objective typically being to maximize the task loss on the unlearning data or to minimize the random label loss. Some methods identify the related parameters and then make appropriate modifications. In-context learning-based methods modify the LLM input prompts to make the LLM refuse to output content related to the unlearning data. Regarding unlearning effectiveness, parameter optimization is typically much more effective than in-context learning.
@@
However, these methods still often poorly maintain the model utility outside the unlearned knowledge, especially in real-world continual settings. The challenges are two-fold: (i): First, in addition to the data that needs to be unlearned, existing unlearning methods also require a large dataset called the retained dataset to maintain the model utility. This retained dataset often consists of the original training dataset (Bourtoule et al., 2021) or a portion of it, but as LLMs are trained on massive datasets (Wang et al., 2024), assuming access to the complete training data is typically unrealistic (Liu et al., 2024). Moreover, as time goes on, the original training data of LLMs may become inaccessible due to expired access authorization, data privacy, and intellectual property protection (Sun et al., 2024). If the retained dataset only contains incomplete training data distribution, the model utility of the missing parts significantly declines after unlearning. Although some studies shrink the range of the retained data to the distribution most susceptible to unlearning, this distribution itself is hard to characterize and its data may be limited due to intrinsic rarity and privacy protection (Chang et al., 2024; Huang et al., 2024). (ii): The second challenge is that existing LLM unlearning methods only consider single operations and cannot perform effective continual unlearning. LLM unlearning is often not a one-off operation but a continual process, as unlearning requests continuously emerge in the real world (Liu et al., 2024). As the number of unlearning operations increases, the aforementioned decline in model utility will also have a cumulative effect, even with the retained dataset, meaning that the model’s general capabilities will significantly decrease Gu et al. (2024); Gupta et al. (2024) over time.
@@
In this work, to achieve more effective continual unlearning for LLMs, we propose the O3 framework, which can balance unlearning effectiveness and model utility preservation in continuous scenarios without using any retained data. At a high level, the O3 framework mainly includes an Orthogonal Low-rank adapter (LoRA) (Hu et al., 2021) for continuously unlearning requested data and an Out-Of-Distribution (OOD) detection module to assess the similarity between input data and unlearning data. Specifically, the orthogonal LoRA in O3 enables the disentanglement of parameter space across different unlearning requests, ensuring that the unlearning effectiveness of different requests does not interfere with each other. Then the OOD detector in O3 is trained with a novel contrastive entropy loss as its backbone and supplemented with a glocal-aware scoring mechanism. The O3 framework can balance unlearning and utility because it smartly leverages the data similarity determined by the OOD detector to decide whether and to what extent to load the unlearning LoRA during inference. In summary, the main contributions of this work include:
  • We study the underexplored problem of LLM continual unlearning and tackle the challenge of balancing unlearning effectiveness and model utility preservation when LLM faces the continuous arrival of unlearning requests, without using any retained data.
@@
  • We propose a novel O3 framework that includes an orthogonal unlearning LoRA and an OOD detector. The orthogonal design of LoRA prevents interference among different unlearning requests, achieving better unlearning effectiveness in continuous scenarios. The OOD detector measures the similarity between input and unlearning data, allowing O3 to decide whether and to what extent to load the unlearning LoRA during inference.
@@
  • We conduct extensive experiments on multiple benchmark tasks that comprehensively test the LLM continual unlearning on discriminative, generative, and reasoning tasks. The experiment results demonstrate that O3 consistently achieves the best balance between unlearning effectiveness and utility preservation without using any retained data, compared with many state-of-the-art baseline methods when facing continuous unlearning requests.
2 Preliminary
Language Model Unlearning. Machine unlearning (Bourtoule et al., 2021) is proposed to protect data privacy and ensure authorized usage (Liu et al., 2024; Zhang et al., 2023). There have been approaches to achieving unlearning through parameter optimization (Jang et al., 2023; Golatkar et al., 2020; Yao et al., 2023; Eldan & Russinovich, 2023; Zhang et al., 2024; Jia et al., 2024; Meng et al., 2022; Yu et al., 2023; Wu et al., 2023; Li et al., 2024), and in-context learning (Thaker et al., 2024; Pawelczyk et al., 2024). The optimization-based unlearning is to employ GradAsc (Golatkar et al., 2020; Yao et al., 2023) on the unlearned data. The following approaches, like PO (Eldan & Russinovich, 2023; Zhang et al., 2024; Jia et al., 2024), notice that unconstrained GradAsc hurts the model’s utility, thus crafting task labels through shuffling or rejection. Yu et al. (2023) localizes the model parameters related to unlearning data and updates them through merging or subtracting (Ding et al., 2023). The in-context learning-based methods adjust input prompts to reject unwanted content generation. Although these approaches can achieve unlearning in certain cases, they neglect that unlearning requests always appear continuously (Liu et al., 2024). The challenge is that the accumulative impact of conducting unlearning sequentially induces catastrophic model utility loss (Gu et al., 2024; Gupta et al., 2024). Besides, their success in a single unlearning operation relies on sufficient retained data with accurate labels, which are scarce in real-world LLM applications. Our O3 addresses the above challenges via orthogonal unlearning LoRA and a designed OOD detector.
Out-of-distribution Detection. Unlike traditional OOD detection based on one-class SVM (Erfani et al., 2016), random forest (Primartha & Tama, 2017), and Gaussian mixture modeling (Laxhammar et al., 2009), deep learning-based OOD detection (Yang et al., 2021) has become mainstream, focusing on the classification task. These approaches assume the availability of category-labeled ID data. However, only a few studies explore unsupervised OOD detection in language models with only unlabeled ID data. DAGMM (Zong et al., 2018) uses a deep auto-encoder to generate low-dimensional representations to estimate OOD scores, only working on multivariate time-series data. Xu et al. (2021) leverage the feature extracted from a pre-trained language model and then fit an OC-SVM for detection. More recent related works search ways like ensemble learning (Zhou et al., 2023), pseudo labeling (Lang et al., 2022), outlier exposure (Cao et al., 2024), and prefix tuning (Ouyang et al., 2023) to achieve OOD text intent detection. Although these works can achieve OOD detection over non-classification tasks, their effectiveness still relies on labels of specific tasks. Our proposed O3 does not require any types of labels.
Problem Definition. In the context of LLM, we consider the prevalent causal language models, which take textual token sequences with variable lengths as the input. As an autoregressive model, the target LLM $M_{\Theta}$ parameterized on $\Theta$ calculates the probability of each token in a text with the previous tokens. For the problem of unlearning, we suppose a sequence of continually arriving unlearning requests with $N^{\mathrm{U},t}$ data samples, i.e., $\{\mathcal{D}^{\mathrm{U},t}\}_{t=1}^{T}$ where, for $t$-th request, $\mathcal{D}^{\mathrm{U},t}=\{\bm{x}_{i}\|\bm{x}_{i}\sim\mathcal{P}_{\mathcal{X}}^{\mathrm{U},t}\}_{i=1}^{N^{\mathrm{U},t}}$ and $\mathcal{P}_{\mathcal{X}}^{\mathrm{U},t}$ is the input marginal distribution. $T$ is an index of the latest request. Each unlearning request corresponds to a task with the training data characterized by the input $\mathcal{P}^{t}_{\mathcal{X}}$ and label distributions $\mathcal{P}^{t}_{\mathcal{Y}}$. Traditional unlearning approaches always assume a retained dataset drawn from a distribution $\mathcal{P}^{\mathrm{R},t}_{\mathcal{X}}$ disjoint from the unlearning one $\mathcal{P}^{\mathrm{U},t}_{\mathcal{X}}$, to retain the model performance on the original training distribution. Then, on one hand, the straightforward objective of continual unlearning is:
$\sum_{t=1}^{T}\min_{\Theta^{t}}\mathbf{I}(M_{\bm{x}\sim\mathcal{P}_{\mathcal{X}}^{\mathrm{U},t}}(\bm{x},\Theta);M^{t}_{\bm{x}\sim\mathcal{P}_{\mathcal{X}}^{\mathrm{U},t}}(\bm{x},\Theta^{t})),\,\sum_{t=1}^{T}\max_{\Theta^{t}}\mathbf{I}(M_{\bm{x}\sim\mathcal{P}_{\mathcal{X}}^{\mathrm{R},t}}(\bm{x},\Theta);M^{t}_{\bm{x}\sim\mathcal{P}_{\mathcal{X}}^{\mathrm{R},t}}(\bm{x},\Theta^{t})),$ (1)